# SpikeStereoNet: A Brain-Inspired Framework for Stereo Depth Estimation from Spike Streams

### v1

@@ 4.4 Ablation Study
To evaluate the effectiveness of each component in our framework and analyze their contributions to the overall performance, we conduct a series of ablation studies in the synthetic dataset.
@@
Table 2: Ablation studies of our proposed modules. All experiments are conducted on the synthetic dataset.
[…]
Removing RC and FFC module 12.07 6.46 1.29
Removing RC module 11.05 4.48 0.83
Removing FFC module 5.86 3.63 0.68
Full model 4.52 2.61 0.44
Removing Reg module 6.38 2.85 0.51
Removing GN module 7.33 2.97 0.57
Removing Reg and GN module 7.77 3.41 0.61
@@ 5 Conclusion
Appendix A Technical Appendices and Supplementary Material
The supplementary materials provide additional implementation details and extended experimental results for the SpikeStereoNet framework introduced in the main paper. It is organized into four distinct sections: Theory Analysis and Method in Appendix  A.1, Details of the Synthetic Dataset in Appendix  A.2, Details of the Real Dataset in Appendix  A.3, More Experimental Results in Appendix  A.4 and Broader Impacts in Appendix  A.5.
A.1 Theory Analysis and Method
A.1.1 Problem Statement
Given a continuous binary spike stream $S_{N}^{t_{1}}\in\{0,1\}^{H\times W\times N}$ with a spatio-temporal resolution of $H\times W\times N$, centered on the timestamp $t_{1}$, we divide it into three non-overlapping sub-streams $S_{N}^{t_{0}}$, $S_{N}^{t_{1}}$, and $S_{N}^{t_{2}}$, centered on timestamps $t_{0}$, $t_{1}$, and $t_{2}$, respectively. The goal is to estimate the disparity map corresponding to $t_{1}$. Here, $S_{N}^{t_{1}}$ provides the essential spatial structure for estimating the disparity, while $S_{N}^{t_{0}}$ and $S_{N}^{t_{2}}$ provide temporal cues to improve consistency and robustness. Letting $S_{n}\in\{0,1\}^{W\times H}$ denote the $n$-th spike frame, we predict the disparity at timestamp $t_{1}$ from the full spike stream $\{S_{n}\},n=0,\dots,k$, leveraging both spatial features and temporal continuity.
A.1.2 Noise Simulation
Noise in spiking cameras originates mainly from the dark current. The accumulation of brightness at pixel $(i,j)$ over time $t$ can be expressed as:
[…]
where $A(i,j,t)$ represents the integrated brightness, $t^{\text{pre}}_{i,j}$ is the previous firing time for pixel $(i,j)$, and $I_{\text{dark}}$ denotes the random noise component caused by dark current. In our synthetic dataset, the dark currents in the circuits introduce thermal noise, which is the type of noise modeled in this work.
A.1.3 Video-to-Spike Preprocessing Pipeline
@@
To enhance temporal resolution, raw frames from dynamic synthetic datasets are processed using a pre-trained video frame interpolation model. The architecture comprises a hierarchical feature extractor, a MultiScaleFlow module for multi-resolution optical flow estimation, and a U-Net for residual refinement through bidirectional flow guidance and mask fusion.
For temporal expansion, we applied a frame rate upsampling factor of ×$\times 50$ to the synthetic dataset. The output is represented as a 4D tensor of shape $[T,H,W,C]$, where $T$ is the temporal length, $H\times W$ is the spatial resolution, and $C=3$ denotes RGB channels.
@@
    spike $\text{spike\ matrix}[t,x,y]=\begin{cases}1&\text{if }V_{t}(x,y)\geq\theta\\ 0&\text{otherwise}\end{cases}$
@@
    Repeat until all frames processed.
Spike streams are stored as binary tensors of shape $[T,H,W]$ using the StackToSpike function with configurable noise $I_{\text{noise}}$ and threshold $\theta$. For compact storage and compatibility for our framework, the SpikeToRaw module compresses spikes (8 per byte) into .dat format, enabling lossless reconstruction during inference.
The proposed two-stage preprocessing pipeline effectively bridges conventional images and neuromorphic vision processing. By combining deep learning-based frame interpolation with bio-inspired spike encoding, we achieve:
  • Temporal Super-Resolution: Neural interpolation extends temporal sampling density by 50× through multi-scale optical flow and attention mechanisms, preserving physical consistency in dynamic scenes.
  • Biologically Plausible Encoding: The temporal integration algorithm emulates retinal neuron dynamics, converting intensity variations into sparse spike events with adaptive threshold control.
@@
To assess the fidelity of the dataset conversion, we utilize the Texture From Interval (TFI) algorithm of the SpikeCV [76] library to reconstruct grayscale images from spike streams with dimensions $[T,H,W]$. This method exploits the spatiotemporal sparsity and encoding capacity of spike streams to approximate the textural structure of conventional images.
The core idea behind TFI is that the temporal interval between adjacent spikes encodes local texture intensity: shorter intervals correspond to higher pixel activity and brighter intensity. The algorithm identifies the two nearest spike timestamps within a bounded temporal window (±$\pm\Delta t$) for each pixel and computes the pixel-wise grayscale value based on their interval duration.
A.1.4 Adaptive Leaky Integrate-and-Fire Neuron Model
@@
In this work, we adopt a simplified formulation. The membrane potential of each neuron is updated using the Exponential Euler Algorithm [77]. The firing threshold $\bm{v_{t}^{th}}$ increases by a fixed value $\bm{{\beta}_{t}}$ following each spike, while the peak voltage $\bm{v_{t}^{\textbf{th}}}=\bm{\beta_{t}}\cdot v_{\text{peak}}$ remains constant. At a discrete time step of $\Delta t=1$ ms, the neuronal membrane potential update rule is defined as:
@@
A.1.5 Applying BPTT to RSNN
@@
the spiking function of sigmoid gradient is used in backpropagation. The gradient is defined by:
@@
where $\alpha$ is a parameter that controls the slope of the surrogate derivative. Unless mentioned otherwise we set $\alpha=4$.
A.1.6 Detail of Loss Function
In our training framework, the total loss $L$ is composed of three distinct components, each serving a specific purpose in optimizing the performance of model. Specifically, the total loss is defined as
@@ A.1.6 Detail of Loss Function
where $L_{\text{stereo}}$ represents the stereo loss, $L_{\text{rate\_reg}}$ denotes the firing rate regularization loss, and $L_{\text{v\_reg}}$ corresponds to the velocity regularization loss. The parameters $\lambda_{f}$ and $\lambda_{v}$ are regularization coefficients that control the influence of the respective regularization terms. The stereo loss $L_{\text{stereo}}$ is designed to measure the discrepancy between the ground-truth disparity $d_{\text{gt}}$ and the predicted disparity $d_{t}$ over time $t$. It is formulated as follows:
$L_{\text{stereo}}=\sum_{t=1}^{T}\eta^{T-t}||d_{\text{gt}}-d_{t}||_{1},$ (10)
@@
where $r_{i}$ is the average firing rate of neuron $i$ over time $T$, this term encourages the model to maintain consistent firing rates across neurons, thereby enhancing stability and preventing over-activation.
The voltage regularization loss $L_{\text{v\_reg}}$ is introduced to regularize the temporal dynamics of the model by penalizing high velocities of neurons. It is expressed as:
@@
Overall, the combination of these loss components allows for a comprehensive optimization strategy that balances accuracy, firing rate stability, and temporal smoothness, thereby enhancing the robustness and performance of the model in stereo vision tasks.
A.2 Details of the Synthetic Dataset
A conventional active stereo depth system typically comprises an infrared (IR) projector, a pair of IR stereo cameras (left and right) and a color camera. Depth measurement is achieved by projecting a dense IR dot pattern onto the scene, which is subsequently captured by the stereo IR cameras from different viewpoints. Depth values are then estimated by applying a stereo matching algorithm to the captured image pair, leveraging the disparity between them. We use the method from [73, 74] replicates this process, containing light pattern projection, capture, and stereo matching. The simulator is implemented based on the Blender [72] rendering engine.
A.2.1 Random Scenes
For each scene, we generate the corresponding ’.meta’ files to store structured metadata information. Additionally, for each object instance, a high-quality ’.obj’ mesh file is created to represent its geometry. To simulate realistic camera motion, a continuous trajectory is generated by sampling camera poses uniformly from a virtual sphere with a fixed radius of 0.6 meters. The representative camera viewpoints are separated by a constant angular increase of 3 °, ensuring smooth transitions along the trajectory. The objects are rotated accordingly to maintain consistent spatial alignment. Furthermore, a uniform background and consistent lighting configuration are applied across all scenes to reduce domain-specific biases and enhance rendering consistency. To further bridge the domain gap between synthetic and real-world data, we adopt domain randomization strategies by randomizing object textures, material properties (from specular, transparent, to diffuse), object layouts, floor textures, and lighting conditions aligned with camera poses.
A.2.2 Stereo Matching
@@
where $Z$ is the obtained depth, $d$ is the disparity, $f$ denotes the camera focal length, $B$ is the baseline distance between the stereo cameras, and $c_{x1}-c_{x0}$ represents the horizontal offset between the principal points of the right and left cameras.
A.2.3 Synthetic Dataset
We make use of domain randomization and depth simulation and construct the large-scale synthetic dataset. In total, the dataset consists of two subsets: (1) Training set: Consists of 76,500 spike streams and 38,250 depth maps generated from 150 scenes composed of 88 distinct objects, with randomized specular, transparent, and diffuse materials. (2) Test set: 20,400 spike streams and 10,200 depth maps generated from 40 new scenes containing 88 distinct objects. The scenes in the dataset are shown in Fig. A1 and Fig. A2, and the detailed statistics of dataset are shown in Table A1.
Table A1: Statistics of the training and test set of synthetic dataset.
@@ A.2.3 Synthetic Dataset
Figure A1: Scenes for generating the training set of the synthetic dataset. Selected five scenes, each displayed as a row in which the images correspond to different camera viewpoints of the same scene (partial views).
@@
Figure A2: Scenes for generating the test set of the synthetic dataset. Selected two scenes, each displayed as a row in which the images correspond to different camera viewpoints of the same scene (partial views).
A.3 Details of the Real Dataset
@@
We construct a stereo imaging system using two spike cameras and one RGB-D camera (Kinect) to simultaneously capture left/right spike streams and the corresponding ground-truth depth images. As illustrated in Fig. A3, the two spike cameras are mounted on a high-precision alignment rig to ensure accurate baseline calibration. The spatial resolution of the spike camera is $400\times 250$, and the temporal resolution is 20,000 Hz. The resolution of the RGB-D camera is $1280\times 720$, and the frequency is 30 fps. Compared with LiDAR, the RGB-D sensor provides dense depth annotations within its effective range, avoiding the sparse and incomplete depth output typically associated. To ensure spatial alignment between modalities, intrinsic and extrinsic calibrations are performed between the spike cameras and the depth camera. The captured depth maps are then geometrically warped to the coordinate frame of the left spike camera for accurate correspondence.

### v2

@@ 4.4 Ablation Study
To evaluate the effectiveness of each component in our framework and analyze their contributions to the overall performance, we conducted a series of ablation studies in the synthetic dataset.
@@
Table 2: Ablation studies of our modules. All experiments are conducted on the synthetic dataset.
Setting of experiment bad 2.0 $(\%)\downarrow$ bad 3.0 $(\%)\downarrow$ AvgErr (px) $\downarrow$
w/o RC & FFC modules 12.07 6.46 1.29
w/o RC module 11.05 4.48 0.83
w/o FFC module 5.86 3.63 0.68
Ours (full) 4.13 2.38 0.42
w/o Reg module 6.38 2.85 0.51
w/o GN module 7.33 2.97 0.57
w/o Reg & GN modules 7.77 3.41 0.61
@@
More ablation studies on the effectiveness of the methods can be found in Appendix A.5.
@@ 5 Conclusion
Ethics Statement
We affirm compliance with the Code of Ethics for all stages of this work. Our study does not involve interventions on human or animal subjects. We evaluate and report fairness indicators where applicable, and we avoid training or releasing models that systematically disadvantage protected groups. All code, models, and preprocessing scripts will be released with documentation to support reproducibility while respecting data licenses and usage constraints. We also report compute settings to encourage consideration of environmental impact and resource use.
Reproducibility Statement
To facilitate reproduction of our results, we provide detailed descriptions of our methodology. All experiments were conducted under PyTorch with CUDA, and the exact environment is specified. We specify all model architectures, hyperparameters, training protocols, data-splitting strategies, and evaluation metrics in the main body and Appendix. All experiments were performed with fixed random seeds. We will release our complete dataset, code repository, and demo.
Appendix A Appendix
The appendix provides additional implementation details and extended experimental results for the SpikeStereoNet framework introduced in the main paper. It is organized into six distinct sections: The Use of Large Language Models in Appendix  A.1, Theory Analysis and Method in Appendix  A.2, Details of the Synthetic Dataset in Appendix  A.3, Details of the Real Dataset in Appendix  A.4, More Experimental Results in Appendix  A.5 and Broader Impacts in Appendix  A.6.
A.1 The Use of Large Language Models (LLMs)
During writing, we use large language models only for grammar and style polishing. The LLMs did not generate or rewrite technical content. All statements and citations were verified by the authors. No proprietary data or restricted code was shared with the LLMs. The authors assume full responsibility for the content of the paper.
A.2 Theory Analysis and Method
A.2.1 Problem Statement
Given a continuous binary spike stream $S_{N}^{t_{1}}\in\{0,1\}^{H\times W\times N}$ with a spatio-temporal resolution of $H\times W\times N$, centered on the timestamp $t_{1}$, we divide it into three non-overlapping substreams $S_{N}^{t_{0}}$, $S_{N}^{t_{1}}$, and $S_{N}^{t_{2}}$, centered on timestamps $t_{0}$, $t_{1}$, and $t_{2}$, respectively. The goal is to estimate the disparity map corresponding to $t_{1}$. Here, $S_{N}^{t_{1}}$ provides the essential spatial structure for estimating the disparity, while $S_{N}^{t_{0}}$ and $S_{N}^{t_{2}}$ provide temporal cues to improve consistency and robustness. Letting $S_{n}\in\{0,1\}^{W\times H}$ denote the $n$-th spike frame, we predict the disparity at timestamp $t_{1}$ from the full spike stream $\{S_{n}\},n=0,\dots,k$, leveraging both spatial features and temporal continuity.
A.2.2 Noise Simulation
The noise in spiking cameras comes mainly from the dark current. The accumulation of brightness in the pixel $(i,j)$ over time $t$ can be expressed as:
$A(i,j,t)=\int_{t^{\text{pre}}_{i,j}}^{t}\left(I(i,j,\tau)+I_{\text{dark}}(i,j,\tau)\right)d\tau,$ (5)
where $A(i,j,t)$ represents the integrated brightness, $t^{\text{pre}}_{i,j}$ is the previous firing time for the pixel $(i,j)$, and $I_{\text{dark}}$ denotes the random noise component caused by the dark current. In our synthetic dataset, the dark currents in the circuits introduce thermal noise, which is the type of noise modeled in this work.
A.2.3 Video-to-Spike Preprocessing Pipeline
@@
To enhance temporal resolution, raw frames from dynamic synthetic datasets are processed using a pre-trained video frame interpolation model EMA-VFI (Zhang et al., 2023a). The architecture comprises a hybrid CNN and transformer framework, and uses correlation information hidden within the attention map to simultaneously enhance the appearance information and model motion.
For temporal expansion, we applied a frame rate upsampling factor of $\times 50$ to the synthetic dataset. The output is represented as a 4D tensor of shape $[T,H,W,C]$, where $T$ is the temporal length, $H\times W$ is the spatial resolution and $C=3$ denotes RGB channels.
@@
    $\text{spike\ matrix}[t,x,y]=\begin{cases}1&\text{if }V_{t}(x,y)\geq\theta\\ 0&\text{otherwise}\end{cases}$
@@
    Repeat until all frames are processed.
Spike streams are stored as binary tensors of shape $[T,H,W]$ using the StackToSpike function with configurable noise $I_{\text{noise}}$ and threshold $\theta$. For compact storage and compatibility for our framework, the SpikeToRaw module compresses spikes (8 per byte) in the .dat format, allowing lossless reconstruction during inference.
The proposed two-stage preprocessing pipeline effectively bridges conventional images and neuromorphic vision processing. By combining deep learning-based frame interpolation with bio-inspired spike encoding, we achieve the following:
  • Temporal Super-Resolution: Neural interpolation extends the temporal sampling density by $50\times$ through multi-scale optical flow and attention mechanisms, preserving physical consistency in dynamic scenes.
  • Biologically Plausible Encoding: The temporal integration algorithm emulates the neuron dynamics of the retina, converting intensity variations into sparse spike events with adaptive threshold control.
@@
To assess the fidelity of the conversion of the dataset, we use the Texture From Interval (TFI) algorithm of the SpikeCV (Zheng et al., 2023) library to reconstruct grayscale images from spike streams with dimensions $[T,H,W]$. This method exploits the spatiotemporal sparsity and encoding capacity of spike streams to approximate the textural structure of conventional images.
The core idea behind TFI is that the temporal interval between adjacent spikes encodes local texture intensity: shorter intervals correspond to higher pixel activity and brighter intensity. The algorithm identifies the two nearest spike timestamps within a bounded temporal window ($\pm\Delta t$) for each pixel and computes the grayscale of the value pixel based on their interval duration.
A.2.4 Adaptive Leaky Integrate-and-Fire Neuron Model
@@
In this work, we adopt a simplified formulation. The membrane potential of each neuron is updated using the Exponential Euler Algorithm (Cachia, 2004). The firing threshold $\bm{v_{t}^{th}}$ increases by a fixed value $\bm{{\beta}_{t}}$ following each spike, while the peak voltage $\bm{v_{t}^{\textbf{th}}}=\bm{\beta_{t}}\cdot v_{\text{peak}}$ remains constant. At a discrete time step of $\Delta t=1$ ms, the neuronal membrane potential update rule is defined as:
@@
A.2.5 Applying BPTT to RSNN
@@
the spiking function of the sigmoid gradient is used in backpropagation. The gradient is defined by:
@@
where $\alpha$ is a parameter that controls the slope of the surrogate derivative. Unless otherwise mentioned, we set $\alpha=4$.
A.2.6 Detail of Loss Function
In our training framework, the total loss $L$ is composed of three distinct components, each serving a specific purpose in optimizing the performance of the model. Specifically, the total loss is defined as
@@ A.2.6 Detail of Loss Function
where $L_{\text{stereo}}$ represents stereo loss, $L_{\text{rate\_reg}}$ denotes the firing rate regularization loss, and $L_{\text{v\_reg}}$ corresponds to the velocity regularization loss. The parameters $\lambda_{f}$ and $\lambda_{v}$ are regularization coefficients that control the influence of the respective regularization terms. Stereo loss $L_{\text{stereo}}$ is designed to measure the discrepancy between the ground-truth disparity $d_{\text{gt}}$ and the predicted disparity $d_{t}$ over time $t$. It is formulated as follows:
[…]
@@
where $r_{i}$ is the average firing rate of the neuron $i$ over time $T$, this term encourages the model to maintain consistent firing rates across neurons, thereby enhancing stability and preventing over-activation.
The voltage regularization loss $L_{\text{v\_reg}}$ is introduced to regularize the temporal dynamics of the model by penalizing the high velocities of the neurons. It is expressed as
@@
In general, the combination of these loss components allows for a comprehensive optimization strategy that balances accuracy, firing rate stability, and temporal smoothness, thereby enhancing the robustness and performance of the model in stereo vision tasks.
A.3 Details of the Synthetic Dataset
A conventional active stereo depth system typically comprises an infrared (IR) projector, a pair of IR stereo cameras (left and right) and a color camera. Depth measurement is achieved by projecting a dense IR dot pattern onto the scene, which is subsequently captured by the stereo IR cameras from different viewpoints. The depth values are then estimated by applying a stereo matching algorithm to the captured image pair, leveraging the disparity between them. We use the method from (Dai et al., 2022; Zhang et al., 2024a) that replicates this process, containing light pattern projection, capture, and stereo matching. The simulator is implemented based on the Blender (Community, 2018) rendering engine.
A.3.1 Random Scenes
For each scene, we generate the corresponding ‘.meta’ files to store structured metadata information. Additionally, for each object instance, a high-quality ‘.obj’ mesh file is created to represent its geometry. To simulate realistic camera motion, a continuous trajectory is generated by sampling camera poses uniformly from a virtual sphere with a fixed radius of 0.6 meters. The representative camera viewpoints are separated by a constant angular increase of 3 °, ensuring smooth transitions along the trajectory. The objects are rotated accordingly to maintain consistent spatial alignment. Furthermore, a uniform background and consistent lighting configuration are applied across all scenes to reduce domain-specific biases and enhance rendering consistency. To further bridge the domain gap between synthetic and real-world data, we adopt domain randomization strategies by randomizing object textures, material properties (from specular, transparent, to diffuse), object layouts, floor textures, and lighting conditions aligned with camera poses.
A.3.2 Stereo Matching
@@
where $Z$ is the depth obtained, $d$ is the disparity, $f$ denotes the focal length of the camera, $B$ is the baseline distance between the stereo cameras and $c_{x1}-c_{x0}$ represents the horizontal offset between the principal points of the right and left cameras.
A.3.3 Synthetic Dataset
We make use of domain randomization and depth simulation and construct a large-scale synthetic dataset. In total, the dataset consists of two subsets: (1) Training set: Consists of 76,500 spike streams and 38,250 depth maps generated from 150 scenes composed of 88 distinct objects, with randomized specular, transparent, and diffuse materials. (2) Test set: 20,400 spike streams and 10,200 depth maps generated from 40 new scenes containing 88 distinct objects. The scenes in the dataset are shown in Fig. 8 and Fig. 9, and the detailed statistics of dataset are shown in Table 3.
Table 3: Statistics of the training and test set of synthetic dataset.
@@ A.3.3 Synthetic Dataset
Figure 8: Scenes for generating the training set of the synthetic dataset. Selected five scenes, each displayed as a row in which the images correspond to different camera viewpoints of the same scene (partial views).
@@
Figure 9: Scenes for generating the test set of the synthetic dataset. Selected two scenes, each displayed as a row in which the images correspond to different camera viewpoints of the same scene (partial views).
A.4 Details of the Real Dataset
A.4.1 Stereo Camera System
@@
We constructed a stereo imaging system using two spike cameras and one RGB-D camera (Kinect) to simultaneously capture left/right spike streams and the corresponding ground-truth depth images. As illustrated in Fig. 10, the two spike cameras are mounted on a high-precision alignment rig to ensure accurate baseline calibration. The spatial resolution of the spike camera is $400\times 250$, and the temporal resolution is 20,000 Hz. The resolution of the RGB-D camera is $1280\times 720$, and the frequency is 30 FPS. Compared with LiDAR, the RGB-D sensor provides dense depth annotations within its effective range, avoiding the sparse and incomplete depth output typically associated. To ensure spatial alignment between modalities, intrinsic and extrinsic calibrations are performed between the spike cameras and the depth camera. The captured depth maps are then geometrically warped to the coordinate frame of the left spike camera for accurate correspondence.
A.4.2 Stereo System Calibration
For the calibration of spike cameras and Kinect, we used the following methods and programs. When calibrating with spike cameras, it is necessary to capture images of a standard chessboard calibration pattern. Approximately 20 sets of images should be taken from different positions and angles (covering left-right and up-down tilts of 15-45 degrees, various distances, and the entire field of view). During shooting, the calibration board must remain stationary, be fully visible in the field of view of both spike cameras, and both cameras should capture images simultaneously. Sufficient time should be allowed at each position to ensure image clarity. Additionally, when reconstructing images using TFI/TFP (Zheng et al., 2023), the hardware and software settings (including the TFI/TFP parameters) of the two spike cameras must be kept consistent. To reduce noise interference, the process should start with completely static scenes and longer TFI windows, ensuring that the scenes captured by both cameras are identical except for perspective differences. Adequate lighting should be provided, and scenes with clear, and strong edges should be selected to improve calibration success rates. OpenCV’s calibration algorithms are used for stereo rectification (the primary APIs are “cv.stereoRectifyUncalibrate” and “cv.stereoRectify”). The rectification is applied to the TFI/TFP images reconstructed by the spike cameras, and the rectified images serve as input to the binocular stereo algorithm.
A.4.3 Real Dataset Collection